OpenStudy (anonymous):

Solve the following system using the substitution method: 2x + 4y = 5 x = -2y + 6

OpenStudy (satan ):

You have what x is already, just plug that in: \(\Large 2(-2y+6)+4y=5\) Now just solve it.

OpenStudy (anonymous):

is there a solution coordinate? or is it dependent or inconsistant?

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

@alyygirl

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok,what does x equal to?

OpenStudy (anonymous):

its no soluion

OpenStudy (anonymous):

but is it inconsistent then or dependent

OpenStudy (anonymous):

inconsistent

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

re ll no solutions inconsistnet

OpenStudy (anonymous):

well, in many systems we need to know what does x (or y) equal to and just substitute it in another equation.here we have 2 equation: x = -2Y +6 2x + 4y = 5 but you know what is x as the question has given: x = -2y + 6 so we can substiute -2y + 6 in x in this equation: 2x + 4y = 5 so it would be: 2(-2y + 6) + 4y = 5 can u do the rest ;)

OpenStudy (anonymous):

are all

OpenStudy (anonymous):

ok got that one @PFEH.1999

OpenStudy (anonymous):

what about??? Solve the following system using the substitution method: y = 4x - 5 12x - 3y = 15

OpenStudy (anonymous):

and here in this one you have y , it equals to 4x - 5 so substitute y in this equation : 12x - 3y = 15

OpenStudy (anonymous):

this one is just like previous one but here you have y , in that one you had x, you should do the same thing ;)

OpenStudy (anonymous):

12x+-3(4x-5)=15 12x-12x+15=15

OpenStudy (anonymous):

+- what does that mean? :"D

OpenStudy (anonymous):

I put that because you need to distribute the negative

OpenStudy (anonymous):

so @Venny what are the coordinates or is it no solution

OpenStudy (anonymous):

well what did you find when you simplified that alyygirl?

OpenStudy (anonymous):

I got always true so its dependent

OpenStudy (anonymous):

its equal

OpenStudy (anonymous):

Yes itss dependent because its always true

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how about this?? Solve the following system using the substitution method: y = 3x - 9 y = -x - 1

OpenStudy (anonymous):

again you can solve it by others way but if you want sybstitution you should find x from this equation: y = -x - 1

OpenStudy (anonymous):

here x = -y - 1

OpenStudy (anonymous):

you can substitute -y - 1 instead of x here: y = -x -9

OpenStudy (anonymous):

@PFEH.1999 You can also plug the -x-1 into the first equation

OpenStudy (anonymous):

yes we can do this too ;)

OpenStudy (anonymous):

Im saying no solution inconsistent?

OpenStudy (anonymous):

Thats not what I got -x-1=3x-9 add x to both sides -1=4x-9 add 9 to both sides 8=4x x=2 now we have to find y from x, so we plug x back into one of the original equations y=3(2)-9 y=6-9 y=-3 so the ordered pair is (2,-3)

OpenStudy (anonymous):

so its neither dependent or inconsistent

OpenStudy (anonymous):

Yeah, can you answer that ?

OpenStudy (anonymous):

thanks got it!

OpenStudy (anonymous):

no problem